Find the Area i am needing a little assistance with these problems, i have been tryint to figure them out but i can't seem to get the right answer. can someone please help me
1) A rectangle with perimeter 56 and width 13
2) A regular polygon with 12 sides of legnth s and apothem 2
(Give your answer in terms of s)
3) Two similar cones have scale factor 5:2, If the total area of the larger cone is 375, find the total of the smaller cone

Answer by stanbon(75887)

Find the Area i am needing a little assistance with these problems,
1) A rectangle with perimeter 56 and width 13
Perimeter = 2(length+width)
56 = 2(length+13)
28 = length + 13
length = 15
Area = width*length
Area = 13*15 = 195 sq. units
------------------------------------------
2) A regular polygon with 12 sides of length s and apothem 2
(Give your answer in terms of s)
Area = (1/2)*apothem*side*number of sides
Area = (1/2)*2*s*12
Area = 12s square units
--------------
3) Two similar cones have scale factor 5:2, If the total area of the larger cone is 375, find the total of the smaller cone.
-------
5:2 means 5x/2x
If 5x=375
x= 75
Then 2x = 150 sq. units (smaller cone area)
----------
Cheers,
Stan H.